Consider the demand for a good. At price Rs 4, the demand for the good is 25 units. Suppose price of the good increases to Rs 5, and as a result, the demand for the good falls to 20 units. Calculate the price elasticity?


1
Expert's answer
2021-11-08T17:02:04-0500

Given: Original Quantity (Q) =25 units, Fall in Qunatity (ΔQ)=−5

(ΔQ)=-5 units, New Qunatity (Q

1

(Q1 =20 units, Original price (P)= ₹4, Rise in price (ΔP)

(ΔP)= ₹ 1, New price (P

1

)=₹5

(P1)=₹5

price . Elasticity of Demand (E

d

)=ΔQ

ΔP


×P

Q


=−5

1


×4

25


=(−)0.8

(Ed)=ΔQΔP×PQ=-51×425=(-)0.8

Demand is less elastic as E

d

<1

Ed<1. Negative sign indicates the inverse relationship between price and the quantity demanded.